OpenStudy (anonymous):

solve for x: 4x+t/r (the 4x+t is a fraction over r)

OpenStudy (anonymous):

(4x+t)/r This can be rewritten as: (4x+t)/r = 0 Multiply both sides by r. (4x+t) = 0r (4x+t) = 0 Remove parentheses: 4x + t = 0 4x = -t Divide through by 4. x = 1/4t

OpenStudy (anonymous):

Does that make sense?